# Adaptive Knowledge Graphs Enhance Medical Question Answering: Bridging the Gap Between LLMs and Evolving Medical Knowledge

### v1

@@ Abstract
Large Language Models (LLMs) have greatly advanced medical Question Answering (QA) by leveraging vast clinical data and medical literature. However, the rapid evolution of medical knowledge and the labor-intensive process of manually updating domain-specific resources can undermine the reliability of these systems. We address this challenge with Adaptive Medical Graph-RAG (AMG-RAG), a comprehensive framework that automates the construction and continuous updating of Medical Knowledge Graphs (MKGs), integrates Chain-of-Thought (CoT) reasoning, and retrieves current external evidence (e.g., PubMed, WikiSearch). By dynamically linking new findings and complex medical concepts, AMG-RAG not only boosts accuracy but also enhances interpretability for medical queries.
Evaluations on the MEDQA and MEDMCQA benchmarks demonstrate the effectiveness of AMG-RAG, achieving an F1 score of 74.1% on MEDQA and an accuracy of 66.34% on MEDMCQA—surpassing both comparable models and those 10 to 100 times larger. Importantly, these improvements are achieved without increasing computational overhead, underscoring the critical impact of automated knowledge graph generation and external evidence retrieval in delivering up-to-date, trustworthy medical insights.
Adaptive Knowledge Graphs Enhance Medical Question Answering: Bridging the Gap Between LLMs and Evolving Medical Knowledge
Mohammad R. Rezaei1∗, Reza Saadati Fard2, Jayson L. Parker3, Rahul G. Krishnan4,5, Milad Lankarany1 1 Department of Biomedical Engineering, University of Toronto 2 Department of Computer Science, Worcester Polytechnic Institute 3 Department of Biology, University of Toronto Mississauga 4 Department of Computer Science, University of Toronto 5 Vector Institute *mr.rezaei@mail.utoronto.ca
@@ 1 Introduction
Medical knowledge expands at a tremendous pace, with new research findings, clinical guidelines, and treatment protocols emerging constantly. Large Language Models (LLMs) have already demonstrated their value in harnessing this vast and evolving information for medical question answering by processing large corpora of domain-specific literature and data Nazi and Peng (2024); Liu et al. (2023).
Figure 1: Performance vs parameter numbers for medical QA on the MEDQA and MEDMCQA datasets. AMG-RAG achieves an F1 score of 74.1% on MEDQA and an accuracy of 66.34% on MEDMCQA, surpassing both comparable models and those that are 10 to 100 times larger in size. See Tables 1 and 2 for more details.
However, a major challenge lies in ensuring that these models remain factually current and can accurately represent complex relationships among medical concepts Rohanian et al. (2024); Yu et al. (2024). Traditional approaches to mitigating these issues involve the use of knowledge graphs, which offer structured, interconnected representations of medical information and can support more nuanced reasoning Huang et al. (2021). Yet, constructing and maintaining such graphs is labor-intensive, time-consuming, and expensive. These burdens are particularly acute in a domain as dynamic as medicine, where new insights quickly render old information out-of-date Yang et al. (2024). To address this pressing issue, we propose an automated framework for constructing and continuously evolving Knowledge Graphs (KGs) specifically tailored to medical question answering. By leveraging LLMs agents and domain-specific search tools, our method autonomously generates graph Medical Knowledge Graphs (MKGs), enriched with descriptive metadata, confidence scores, and relevance indicators. In doing so, it drastically reduces the manual effort traditionally required to build and update knowledge graphs, while ensuring alignment with the latest medical advances. Unlike traditional Retrieval Augmented Generation (RAG) solutions that rely heavily on vector similarity for retrieval Lewis et al. (2020), our knowledge-graph-based approach provides more sophisticated reasoning capabilities through shared attributes and explicit relationships. This facilitates accurate synthesis of information across diverse medical domains, ranging from drug interactions and clinical trial data to patient histories and treatment guidelines.
A central component of our solution is the integration of these evolving knowledge graphs into a RAG-based pipeline. New and updated graph entities are continuously fed into an LLM question answering module, ensuring that responses draw upon the most up-to-date and contextually relevant medical information Singhal et al. (2022). Building on this dynamic architecture, we introduce an iterative pipeline, AMG-RAG, which combines insights from these automatically maintained graphs with traditional textual retrieval and multi-step chain-of-thought reasoning. By optimizing retrieval through confidence scoring and adaptive graph traversal, AMG-RAG demonstrates significantly improved accuracy and completeness in medical QA Trivedi et al. (2022).
We assess AMG-RAG on the MEDQA and MEDMCQA benchmarks, which are designed to test evidence retrieval, complex reasoning, and multi-choice comprehension in the medical domain. Our model achieves an F1 score of 74.1% on MEDQA and an accuracy of 66.34% on MEDMCQA, outperforming both similarly sized RAG approaches and much larger state-of-the-art models (Fig. 1). Crucially, these gains do not necessitate additional fine-tuning or higher inference costs; instead, they result from seamlessly integrating knowledge graphs and domain-specific search tools. This efficient and scalable approach underscores the value of dynamically evolving knowledge retrieval in medical QA, offering an avenue for enhancing clinical decision-making by delivering reliable, relationally enriched insights Zhou et al. (2023).
@@ 2 Related Work
Medical QA systems are essential for enhancing clinical decision-making, research, and patient care Nazi and Peng (2024); Liu et al. (2023); Rezaei et al. (2024). Over time, the field has evolved with various technological advancements addressing key challenges in processing medical information Singhal et al. (2022). Domain-specific language models such as BioBERT Lee et al. (2020), PubMedBERT Gu et al. (2021), and MedPaLM Singhal et al. (2023) have achieved significant success in biomedical tasks Rohanian et al. (2024). However, these models often struggle with synthesizing complex relationships between medical entities and integrating data from diverse sources, particularly for rare conditions, drug interactions, and comorbidities Zhou et al. (2023); Yu et al. (2024).
To overcome these challenges, RAG frameworks Rezaei and Dieng (2025); Lewis et al. (2020) have enhanced LLMs by integrating external knowledge sources. Systems like MMED-RAG Xia et al. (2024) have extended this paradigm to include multimodal data. The introduction of Chain-of-Thought (CoT) reasoning has further improved QA performance, with IRCoT Trivedi et al. (2022) combining CoT reasoning with RAG for more sophisticated inference. Recent advancements, such as Gemini’s multimodal and long-context reasoning capabilities, have set new benchmarks in MedQA, surpassing GPT-4 in performance Saab et al. (2024). However, these systems often struggle to adapt to novel queries and dynamic data due to their rigid architectures.
KG-based approaches provide another avenue for advancing medical information processing. Systems like KG-Rank Huang et al. (2021) utilize structured knowledge representations and ontologies to enable hierarchical reasoning and inference. By combining knowledge graphs with ranking and re-ranking techniques, these systems enhance the factual accuracy of long-form QA Yang et al. (2024). However, KG-based systems face significant challenges in maintaining scalability and staying current with rapidly evolving biomedical discoveries.
Difference and Importance of AMG-RAG
Our AMG-RAG dynamically constructs relational medical KGs integrated with advanced search capabilities. Unlike traditional static systems, our approach extracts medical terms from queries, enriches them with real-time data, and utilizes LLMs to infer relationships. This dynamic mechanism ensures continuous alignment with emerging medical knowledge, addressing the limitations of static knowledge bases and pre-trained models. By combining dynamic KGs with CoT reasoning and RAG, our framework improves the adaptability and reliability of medical QA systems.
@@ 3 Method
Knowledge Graphs (KGs) offer structured frameworks for evidence-based reasoning. However, traditional KGs struggle to adapt to dynamic and evolving queries, as well as the continuous influx of new research and evidence in the medical domain. To overcome this limitation in the medical Question Answering (QA) pipeline, we propose a novel approach that automatically constructs medical-KGs using a combination of a Large Language Model (LLM) agent and a specialized medical search tool. This Medical Knowledge Graph (MKG) enhances medical QA by tailoring the graph’s construction and querying processes to each specific query. This section outlines the methodologies used to develop the MKG and the Adaptive Medical Graph-RAG (AMG-RAG) pipeline, aimed at enhancing QA systems with advanced capabilities for retrieval, reasoning, and generation.
3.1 Retrieval Augmented Generation (RAG) for QA
Retrieval Augmented Generation (RAG) is a framework designed to enhance QA by integrating relevant external knowledge into the generation process. The framework combines retriever and generator components to ensure responses are grounded in evidence. Below, we outline various approaches within the RAG paradigm:
3.1.1 RAG
In the RAG approach, the retriever fetches a fixed number of relevant documents, $\{{\mathbf{d}}_{1},{\mathbf{d}}_{2},\ldots,{\mathbf{d}}_{n}\}\in\mathbf{D}$, based on the query ${\mathbf{q}}$. Here, $\mathbf{D}$ represents the set of all domain-specific documents utilized. These documents are concatenated and passed directly to a LLM-based text generator, $G$, which produces the answer $\hat{{\mathbf{a}}}$:
@@ 3.1.1 RAG
3.1.2 RAG with Chain-of-Thought (CoT)
@@ 3.1.3 RAG with Search
Figure 2: Model Schema. A) The pipeline for creating the MKG using search tools and an LLM agent. B) An example of the generated MKG in Neo4J, illustrating nodes and relationships derived from search results and contextual information. C) The AMG-RAG pipeline. D) A simplified RAG pipeline.
@@ 3.2 Medical QA with AMG-RAG
Below, we introduce the proposed AMG-RAG pipeline and outline the process for constructing the MKG, which is detailed in the following section.
The AMG-RAG pipeline consists of the following steps:
  1. 1.
    Question Parsing: Extract medical terms $\{{\mathbf{n}}_{1},{\mathbf{n}}_{2},\ldots,{\mathbf{n}}_{m}\}$ from the user query ${\mathbf{q}}$ using an LLM agent:
    The LLM agent is instructed to extract the medical terms $\{{\mathbf{n}}_{i}\}_{i=1}^{m}$ associated with the query ${\mathbf{q}}$.
  2. 2.
    Node Exploration: For each term ${\mathbf{n}}_{i}$, query the KG to retrieve relevant information while applying a confidence threshold. This threshold determines the minimum level of relationship confidence required to include information from the KG, filtering out relationships with confidence scores below the specified threshold. Details on the calculation of confidence scores can be found in Appendix A. Nodes and their children are examined iteratively, with parent confidence scores, $s({\mathbf{n}}_{i})$, multiplied by their relationship scores, $s({\mathbf{r}}_{ij})$, to compute child confidence:
    Both Breadth-First Search (BFS) and Depth-First Search (DFS) strategies can be employed to explore child nodes. BFS prioritizes covering all immediate neighbors at the current depth before moving deeper, ensuring comprehensive breadth-wise exploration. In contrast, DFS delves deeply along one branch before backtracking, enabling a more targeted depth-first traversal. The exploration process continues until either the cumulative confidence meets or exceeds a threshold $\tau$, or the maximum document limit $M$ is reached.
  3. 3.
    Chain of Thought Generation: Generate a reasoning trace ${\mathbf{c}}_{i}$ for each entity ${\mathbf{n}}_{i}$ using an LLM, integrating information from nodes and their relationships:
  4. 4.
    Answer Synthesis: Aggregate reasoning traces $\{{\mathbf{c}}_{1},{\mathbf{c}}_{2},\ldots,{\mathbf{c}}_{m}\}$ and pass them to a final answer generator, which produces the output $\hat{{\mathbf{a}}}$ along with an overall confidence score:
This pipeline ensures that answers are comprehensive and grounded in the KG, with confidence scores providing interpretability and reliability.
@@ 3.3 Dynamic Generation of the Medical Knowledge Graph
The construction of the MKG for QA represents a critical step toward enabling structured reasoning in our AMG-RAG. This approach extracts key entities and their interconnections from user queries, enriching them with information retrieved through external search tools. By organizing information into the MKG, we enable efficient, interpretable, and evidence-based QA. The methodology is as follows:
3.3.1 Node Extraction
Medical terms are identified within the user query ${\mathbf{q}}$ using an LLM agent named Medical Entity Recognizer (MER). These terms are treated as nodes, $\{{\mathbf{n}}_{1},{\mathbf{n}}_{2},\ldots,{\mathbf{n}}_{m}\}$, in the KG. For each extracted term, a search tool (e.g., PubMed, or a specialized medical search engine) retrieves detailed descriptions ${\mathbf{d}}({\mathbf{n}}_{i})$, providing context for each node:
@@ 3.3.1 Node Extraction
The retrieved descriptions form the foundational data for each node in the MKG, ensuring an accurate representation of medical terms and their attributes.
3.3.2 Relationship Inference
An LLM agent extracts relationships between nodes based on their descriptions and retrieved documents. The LLM analyzes pairs of nodes $({\mathbf{n}}_{i},{\mathbf{n}}_{j})$ to determine potential relationships ${\mathbf{r}}_{ij}$ and their nature:
@@ 3.3.2 Relationship Inference
The agent generates a summary, infers the relationship type (e.g., causation, association), and assigns a confidence score. This process results in a KG rich in structure and semantics.
3.3.3 Knowledge Graph Construction
The nodes, descriptions, relationships, and confidence scores are integrated into the KG structure. The resulting graph supports medical QA by:
  • Highlighting key medical concepts and their interrelations.
  • Enabling efficient retrieval and reasoning over medical knowledge.
  • Providing confidence metrics for each established relationship in the graph, serving as a source of reliability.
@@ 4.1 Medical Knowledge Graph
The KG was dynamically constructed by integrating search items, contextual information, and relationships derived from textbooks and search queries from the PubMed engine for each question in the dataset. This data was processed and stored in a Neo4j database. The key features of the knowledge graph include:
@@
    Dynamic Node and Relationship Creation: Nodes are dynamically generated for search items, and relationships between these nodes are established based on their relevance and predefined relationship types.
@@
    Bidirectional Relationships: To ensure a comprehensive representation, the graph includes both forward and reverse relationships between nodes, enhancing its utility for diverse queries.
@@
    Relevance Scoring: Each relationship is enriched with descriptive annotations and a confidence score, quantifying the strength of the association and aiding in prioritizing relevant connections.
@@
    Summarization: Concise summaries for each search item are included, derived from contextual data. A confidence score accompanies each summary to indicate its reliability.
@@
    Integration with Neo4j: The entire graph is stored in a Neo4j database, leveraging its graph-based query capabilities for efficient analysis and retrieval.
A snapshot of a portion of the knowledge graph is shown in Figure 2.B, illustrating its structure and relationships.
This MKG serves as the foundational information source for the AMG-RAG framework during the inference phase. The evaluation of MKG confirmed its robustness and reliability, with experts LLMs such as GPT-4 achieving high precision (e.g.  9/10). These results underscore the effectiveness of MKG in supporting medical reasoning and decision-making, as detailed in Appendix B.

### v2

# Agentic Medical Knowledge Graphs Enhance Medical Question Answering: Bridging the Gap Between LLMs and Evolving Medical Knowledge

@@ Abstract
Large Language Models (LLMs) have greatly advanced medical Question Answering (QA) by leveraging vast clinical data and medical literature. However, the rapid evolution of medical knowledge and the labor-intensive process of manually updating domain-specific resources can undermine the reliability of these systems. We address this challenge with Agentic Medical Graph-RAG (AMG-RAG), a comprehensive framework that automates the construction and continuous updating of Medical Knowledge Graphs (MKGs), integrates reasoning, and retrieves current external evidence from the MKGs for medical QA. Evaluations on the MEDQA and MEDMCQA benchmarks demonstrate the effectiveness of AMG-RAG, achieving an F1 score of 74.1% on MEDQA and an accuracy of 66.34% on MEDMCQA—surpassing both comparable models and those 10 to 100 times larger. By dynamically linking new findings and complex medical concepts, AMG-RAG not only boosts accuracy but also enhances interpretability for medical queries, which has a critical impact on delivering up-to-date, trustworthy medical insights (GitHub: https://github.com/MrRezaeiUofT/AMG-RAG).
Agentic Medical Knowledge Graphs Enhance Medical Question Answering: Bridging the Gap Between LLMs and Evolving Medical Knowledge
Mohammad R. Rezaei1,2∗, Reza Saadati Fard3, Rahul G. Krishnan1,2, Milad Lankarany1 1 University of Toronto 2 Vector Institute 3 Worcester Polytechnic Institute *mr.rezaei@mail.utoronto.ca
@@ 1 Introduction
Medical knowledge is growing at an unprecedented rate: every day brings new research findings, revised clinical guidelines, and updated treatment protocols. Recent work shows that Large Language Models (LLMs) can already harness this ever-expanding corpus for medical Question Answering (QA) Nazi and Peng (2024); Liu et al. (2023).
Figure 1: Performance versus parameter count on the MEDQA and MEDMCQA benchmarks. Our system, AMG-RAG, attains an F1 of 74.1 % on MEDQA and an accuracy of 66.34 % on MEDMCQA, outperforming models that contain 10–100× more parameters. See Tables 1 and 2 for details.
Despite their promise, LLMs face two persistent challenges. First, they must remain factually current in a field where knowledge can become obsolete almost overnight Rohanian et al. (2024); Yu et al. (2024). Second, they must correctly model the intricate relationships among medical entities. Knowledge Graph (KG) provide a structured and interconnected view that supports nuanced reasoning Huang et al. (2021), yet creating and maintaining them by hand is costly—especially in medicine, where new insights rapidly invalidate older facts Yang et al. (2024).
We introduce an automated framework that constructs and continuously refines Medical Knowledge Graphs (MKGs) for QA. Our LLM -driven agents, assisted by domain-specific search tools, generate graph entities enriched with metadata, confidence scores, and relevance indicators. This automation sharply reduces manual curation while keeping the graph aligned with the latest discoveries. In contrast to Retrieval Augmented Generation (RAG) systems that rely solely on vector similarity Lewis et al. (2020), our graph-centric retrieval leverages explicit relationships to synthesise information across domains such as drug interactions, clinical trials, patient histories, and guidelines.
The AMG-RAG combines dynamically synthesised MKG with multi-step reasoning, guided by confidence scores and adaptive traversal strategies Trivedi et al. (2022). This design yields more accurate and complete answers without incurring additional fine-tuning or inference costs. On the MEDQA and MEDMCQA benchmarks—both of which test evidence retrieval, complex reasoning, and multiple-choice comprehension—AMG-RAG achieves an F1 of 74.1% and an accuracy of 66.34%, respectively (Fig. 1). These results surpass those of similarly sized RAG approaches and even much larger state-of-the-art models, underscoring the benefits of a dynamically evolving MKG for medical QA. Our findings highlight the potential of automated, relationally enriched knowledge retrieval to enhance clinical decision-making by delivering timely and trustworthy insights Zhou et al. (2023).
Contributions.
Our contributions are threefold:
  1. 1.
    We developed a autonomous search and graph-building process powered by specialized LLM agents that continuously generate and refine MKGs through integrated workflows using search engines and medical textbooks.
  2. 2.
    Our system embeds confidence scoring mechanisms that explicitly model information uncertainty, providing transparent reliability assessments for medical information.
  3. 3.
    We created an adaptive graph traversal system that transcends traditional retrieval methods, enabling dynamic contextualization of medical knowledge.
@@ 2 Related Work
Medical QA has progressed through three complementary lines of research: (i) domain-specific language models, (ii) retrieval-augmented generation, and (iii) knowledge-graph reasoning.
Domain-specific language models.
BioBERT Lee et al. (2020), PubMedBERT Gu et al. (2021), and MedPaLM Singhal et al. (2023) adapt transformer pre-training to biomedical corpora, delivering strong gains on entity recognition, relation extraction, and multiple-choice QA Nazi and Peng (2024); Liu et al. (2023). Yet, even these specialised models struggle to synthesise multi-hop relations (e.g. rare comorbidities or drug–gene interactions) and must be re-trained to absorb new discoveries Rohanian et al. (2024); Yu et al. (2024).
Retrieval-Augmented Generation (RAG).
RAG pipelines couple an LLM with an external evidence retriever, injecting fresh context at inference time Lewis et al. (2020). Vendi-RAG Rezaei and Dieng (2025) and MMED-RAG Xia et al. (2024) extend this idea to biomedical and multimodal sources, respectively. Chain-of-Thought (CoT) prompting further boosts reasoning: IRCoT Trivedi et al. (2022) interleaves iterative retrieval with step-wise justification. Gemini’s long-context model recently pushed MedQA scores beyond GPT-4 Saab et al. (2024). Nevertheless, most RAG systems rely on static vector stores and cannot explain answers in terms of explicit biomedical relations.
Knowledge-graph reasoning.
KG-Rank Huang et al. (2021) and related work such as KG-RAG Sanmartin (2024) harness ontologies to re-rank evidence or enforce logical constraints, improving factual consistency in long-form QA Yang et al. (2024). However, constructing and curating a high-coverage, up-to-date MKG remains labour-intensive, limiting scalability and freshness.
@@ 3 Method
We propose our framework, Agentic Medical Graph-RAG (AMG-RAG), bridges these threads by dynamically generating a confidence-scored Medical Knowledge Graph (MKG) that is tightly coupled to a Retrieval Augmented Generation (RAG) +CoT pipeline. AMG-RAG features autonomous Knowledge Graph (KG) evolution through Large Language Model (LLM) agents extracting entities and relations from live sources with provenance tracking; graph-conditioned retrieval that maps queries onto the MKG to guide evidence selection; and reasoning over structured context where the answer generator utilizes both textual passages and traversed sub-graphs for transparent, multi-hop reasoning.
3.1 Retrieval Augmented Generation (RAG)
RAG is a framework designed to enhance Question Answering (QA) by integrating relevant external knowledge into the generation process.In the RAG approach, the retriever fetches a fixed number of relevant documents, $\{{\mathbf{d}}_{1},{\mathbf{d}}_{2},\ldots,{\mathbf{d}}_{n}\}\in\mathbf{D}$, based on the query ${\mathbf{q}}$. Here, $\mathbf{D}$ represents the set of all domain-specific documents utilized. These documents are concatenated and passed directly to a LLM-based text generator, $G$, which produces the answer $\hat{{\mathbf{a}}}$:
@@ 3.1 Retrieval Augmented Generation (RAG)
RAG with Chain-of-Thought (CoT).
@@
Figure 2: Model Schema. A) The pipeline for creating the MKG using search tools and an LLM agent. B) An example of the generated MKG in Neo4J, illustrating nodes and relationships derived from search results and contextual information. Our model successfully retrieved and utilized recent knowledge to accurately answer a medical question, highlighting the practical benefit of continuously updating the knowledge graph. Furthermore, we extended this evaluation by providing additional examples retrieved by our system using recent publications in Table 11. C) The AMG-RAG pipeline. D) A simplified RAG pipeline.
@@ 3.2 Medical QA with AMG-RAG
The suggested AMG-RAG framework dynamically creates a MKG and incorporates sophisticated reasoning abilities, overcoming the shortcomings of traditional methods. Our system utilizes structured medical knowledge and reasoning, ensuring flexibility to accommodate new data.
The AMG-RAG pipeline begins with question parsing, where an LLM agent extracts medical terms $\{{\mathbf{n}}_{1},{\mathbf{n}}_{2},\ldots,{\mathbf{n}}_{m}\}$ from the user query ${\mathbf{q}}$:
During node exploration, the system queries the KG for each term ${\mathbf{n}}_{i}$, applying a confidence threshold that filters relationships based on their reliability scores. The system propagates confidence through the KG by computing child confidence as:
Our framework supports both breadth-first and depth-first exploration strategies, enabling flexible knowledge traversal based on query characteristics. The exploration continues until either cumulative confidence meets threshold $\tau$ or document limit $M$ is reached, ensuring comprehensive yet focused information gathering.
The chain-of-thought generation phase synthesizes reasoning traces ${\mathbf{c}}_{i}$ for each entity by integrating information from connected nodes:
Finally, answer synthesis aggregates these reasoning traces to produce the final output $\hat{{\mathbf{a}}}$ with an associated confidence score:
This approach ensures that answers are comprehensive, interpretable, and anchored in reliable medical knowledge.
@@ 3.3 Dynamic Generation of the Medical Knowledge Graph
The development of the MKG marks a pivotal advancement within our AMG-RAG framework, facilitating organized reasoning through the use of dynamically synthesized knowledge representations. Unlike traditional static knowledge bases, our method allows the graph structure to adapt continuously in response to new queries and the evidence they uncover. Furthermore, it assigns a confidence score to each edge in the graph, indicating the reliability of each relationship, which is crucial given the uncertainty inherent in medical evidence.
Node Extraction.
Our Medical Entity Recognizer (MER) agent identifies domain-specific terms within user queries, establishing them as foundational nodes $\{{\mathbf{n}}_{1},{\mathbf{n}}_{2},\ldots,{\mathbf{n}}_{m}\}$ in the knowledge graph. For each identified entity, specialized medical search tools retrieve contextual descriptions ${\mathbf{d}}({\mathbf{n}}_{i})$:
@@
These descriptions provide rich semantic context for the knowledge graph, ensuring accurate representation of medical concepts and their attributes.
Relationship Inference.
The power of our approach lies in its ability to dynamically infer relationships between medical entities. An LLM agent analyzes pairs of nodes $({\mathbf{n}}_{i},{\mathbf{n}}_{j})$ to determine potential relationships ${\mathbf{r}}_{ij}$ and quantify their reliability:
@@
This process generates not just connection types (e.g., causation, correlation, contraindication) but also confidence scores that propagate through subsequent reasoning steps, enabling evidence-weighted inference.
Knowledge Graph Construction.
The resulting MKG integrates extracted entities, their descriptions, inferred relationships, and confidence metrics into a cohesive structure. This graph serves as both a repository of medical knowledge and a computational framework for reasoning. By attaching reliability scores to every edge, downstream components can appropriately weight evidence during inference, enhancing both accuracy and explainability. The graph’s dynamic nature allows it to continuously incorporate new information and refine existing connections, addressing the staleness issues that plague traditional knowledge-based systems.
@@ 4.1 Medical Knowledge Graph
To address the challenges of inaccurate knowledge updating—such as those stemming from noisy retrieval results or LLM hallucinations—our AMG-RAG introduces a robust and dynamic approach to MKG construction. This is particularly critical in healthcare applications, where the absence of error detection and correction mechanisms in automated KG generation can compromise system reliability.
The dynamic update mechanism encompasses strategies resilient to errors, defined by the confidence level of the medical information retained within the MKG and among nodes $i$ and $j$, denoted as ${\mathbf{s}}_{ij}$. This approach facilitates monitoring and reduces the spread of erroneous information during the refinement or reasoning stages. These protective measures allow the system to identify and rectify inconsistencies that may arise from external retrieved information.
The MKG is dynamically constructed for each question by integrating search items, contextual information, and relationships extracted from medical textbooks and search tools, including Wikipedia (Wiki-MKG) and PubMed (PubMed-MKG) queries. The ablation in Table 3 demonstrates that the created MKG based on PubMed (PubMed-MKG) is more effective in enhancing the performance of the AMG-RAG. This data is processed and structured within a Neo4j database. Key innovations in the knowledge graph include:
@@
    Dynamic Node and Relationship Creation: Nodes are instantiated based on retrieved entities and search terms, while relationships are constructed using predefined semantic templates aligned with medical ontologies.
@@
    Bidirectional Relationships: The graph includes both forward and reverse relationships between nodes to allow flexible traversal and comprehensive context understanding.
@@
    Confidence-Based Relevance Scoring: Each relationship is enriched with textual annotations and a quantitative confidence score that measures the reliability of the connection. This confidence score enables the system to down-rank or filter out uncertain associations, thereby mitigating the effects of noisy retrievals.
@@
    Summarization with Reliability Indicators: Each search item is paired with a concise summary derived from contextual sources. These summaries are accompanied by confidence scores that indicate their trustworthiness, allowing nuanced uncertainty modeling.
@@
    Thresholding for Quality Control: In our experiments, we applied a confidence threshold of 8 (on a 10-point scale) to retain only high-reliability nodes and edges. This value was empirically found to yield the best results in benchmark performance.
  6. 6.
    Integration with Neo4j: The complete graph is stored in a Neo4j database, leveraging its powerful graph query engine for efficient retrieval and analysis during inference.
A partial visualization of the MKG structure is shown in Figure 2.B. Additional complete examples with retrieved papers are provided in Table 11. This MKG forms the core knowledge source for the AMG-RAG inference pipeline.
As discussed in Appendix B, extensive validation was conducted through both human and machine evaluators. Clinical experts verified the correctness of the knowledge graph, and expert LLMs such as GPT-4 achieved high accuracy (e.g., 9/10) in validating the extracted knowledge. These results underscore the MKG’s ability to support reliable and explainable medical reasoning within AMG-RAG.
The knowledge graph creation process in AMG-RAG operates independently from the QA process, allowing for continuous background updates of the MKG via search tools such as PubMedSearch or WikiSearch. This approach significantly reduces latency during question answering since the system frequently retrieves information from the pre-populated MKG rather than performing new searches. By maintaining an updated MKG, AMG-RAG achieves a balanced minimum dependency on computational resources and search tools during the test phase.
Despite having only 8B parameters, it delivers competitive results compared to much larger models like Med-Gemini (1800B) and GPT-4 (1760B). Even in worst-case scenarios where relevant information is absent from the MKG, the additional search cost is still significantly lower than the resource requirements of much larger models.